A merchant blends tea that sells for $4.25 a pound with tea that sells for $1.95 a pound to produce 80 lb of a mixture that sells for $2.75 a pound.
Then the mixture contains _____ lbs of $4.25/lb tea and ______ lbs of $1.95 tea

Answer by mananth(16949)

A merchant blends tea that sells for $4.25 a pound with tea that sells for $1.95 a pound to produce 80 lb of a mixture that sells for $2.75 a pound.
Then the mixture contains _____ lbs of $4.25/lb tea and ______ lbs of $1.95 tea

Let the $ 4.25 tea mixed be x lbs.
Let the $ 1.95 tea mixed be y lbs.
x+y = 80 lbs
80 lbs f blended tea at $2.75 . Its vale = 80*2.75 =220
The value of $4.25 tea when mixed will be 4.25x
The value of $1.95 tea when mixed will be 1.95y
4.25x+1.95y=220
x+y=80 Multiply this equation by 1.95
1.95x+1.95y=80*1.95 subtract one from the other
2.30x =64
x=64/2.3 = 27.83 lbs.
80-27.83= 52.17 lbs.

M ananth_________________ mananth@hotmail.com